a wallet contains the same number of pennies,nickels and dimes the coins total is $1.44. How many of each type of coins does the wallet contains if the pennies is equals to 1 centavo,nickels is 5 centavo and the dimes is 10 centavo?
Answer by checkley71(8403)

1N+5N+10N=144
16N=144
N=144/16
N=9 COINS OF EACH VALUE.
PROOF
1*9+5*9+10*9=144
9+45+90=144
144=144